OpenStudy (anonymous):

verify identity: cos 4u= 8cos^4 u - 8cos^2 u + 1

OpenStudy (anonymous):

remember- >\[\cos(2u)=\cos^2u-\sin^2u=2\cos^2u-1\] start from the big one the right side \[ 8\cos^4u-8\cos^2u+1=4(2\cos^2u)^2-4(2\cos^2u)+1\\ =2[\cos(2u)+1]^2-4[\cos(2u)+1]+1\\ =2[\cos^2(2u)+2\cos(2u)+1]-4\cos(2u)-4+1\\ =[2\cos^2(2u)]+4\cos(2u)+2-4\cos(2u)-3\\ =[\cos(4u)+1]-1\\ =\cos(4u) \]

OpenStudy (anonymous):

thanksss !